# Picosecond laser test unit for photosensor characterization

### v1

@@ 4.3 Transit time spectrum
A KDE with appropriate bandwidth is used to smooth the distribution and to locate minima, which serve to separate pulse populations. Following [PMT_SmiLomRan2004, TTandCh_SPE_R7081_PMTs_2012], three populations are identified: prepulses (PPs), main pulses, and late pulses (LPs). Their fractions to the complete data set are counted in the KDE-defined intervals, with an optional dark rate correction estimated from the constant pre-PP region if present. Local full width at half maxima (FWHMs) and full width at quarter maxima (FWQMs) of the pulse populations are taken from the KDE by linear interpolation, and uncertainties are propagated. The three populations are described in the following:
@@
    Prepulses occur $\Delta t_{\text{pre-main}}\simeq 12.6$ ns before the main peak and contribute $0.02\,\%$ of events. The $\text{FWHM}_{\text{PP}}=(1.9\pm 0.1)$ ns is comparable to the main component’s $\text{FWHM}_{\text{TTS}}=(2.0\pm 0.1)$ ns. PPs arise when photons pass the photocathode and strike the first dynode directly [PMT_prelate_2000], yielding smaller charge because the first amplification stage is bypassed [TTandCh_SPE_R7081_PMTs_2012]. Following [PMT_SmiLomRan2004], adapting a photocathode-to-first-dynode distance of $\sim 2.5$ cm (time-of-flight $t_{\text{ToF}}\simeq 0.1$ ns) and the measured pre-main delay $\Delta t_{\text{pre-main}}$, the PE drift time for the 9821B is $t_{\text{PE\,drift}}=\Delta t_{\text{pre-main}}+t_{\text{ToF}}\simeq 12.7$ ns.
@@
    Main pulses are dominant ($98.58\,\%$) and occur from 8.5-8.5 ns to $11.9$ ns with $\text{FWHM}_{\text{TTS}}=(2.0\pm 0.1)$ ns, which is close to the value in the 9821B datasheet [9821B_datasheet]. The TTS reflects variations in PE paths/velocities and secondary-electron production, generally increasing with PMT size. Within the main component we distinguish (i) early pulses from scattering that bypasses an amplification stage (slightly lower charge), (ii) regular pulses that mainly define the TTS, and (iii) delayed pulses from weaker acceleration near photocathode edges [TTandCh_SPE_R7081_PMTs_2012].
@@
    Late pulses comprise $1.40\,\%$ of events, start at $11.9$ ns and peak at $27.7$ ns. They originate from (in)elastic backscattering of photoelectrons at the first dynode. The longest delays are $\approx 2\,t_{\text{PE\,drift}}\simeq 25.4$ ns, consistent with but slightly below the observed LP peak-to-main pulse peak separation due to trajectory spread [PMT_SmiLomRan2004, TTandCh_SPE_R7081_PMTs_2012, PMT_Hama12inch_2013]. Elastic backscatter tends to cluster near the LP peak with charge similar to normal events, while inelastic cases arrive earlier with reduced charge [PMT_prelate_2000]. A secondary bump broadens the LP component, giving $\text{FWHM}_{\text{LP}}=(5.4\pm 0.1)$ ns and $\text{FWQM}_{\text{LP}}=(14.5\pm 0.1)$ ns.
Fig 7 presents charge-time correlations using the trace integral (top) and pulse integral (bottom) methods. The trace integral method illustrates that prepulses are absent within the SPE interval. Due to the systematically smaller charge of PPs, integrating the full waveform, i.e., signal plus surrounding noise, masks their presence.
@@ 5 Results and discussion
This section summarizes the performance of eight 9821B and two 9821QB PMTs from ET Enterprises as well as one Hamamatsu R9980 PMT assembly, characterized at SPE intensity over a range of supply voltages. Although the R9980 PMT requires a negative bias, all voltages are quoted by magnitude. All investigated 9821B devices are equipped with B19A basic hardpin bases, MS75/90 mu-metal shields and E638PFP-01 passive encapsulated voltage dividers. Unless stated otherwise, measurements refer to room temperature and short signal cables in the light‑tight enclosure. Additional details are available in [PhD_MRStock2025].
@@ 5.1 Gain
Figure 8 (top) shows the gain-voltage curves obtained from the trace charge analysis. For each device, an exponential model, $G(U)=G_{0}\exp{\left(U/U_{0}\right)}$, where $G_{0}$ and $U_{0}$ are two fit parameters, describes the data well. Best‑fit curves follow the general scaling suggested by the manufacturer’s divider A benchmark. The parameters $G_{0}$ and $U_{0}$ vary tube‑to‑tube, emphasizing device‑specific behavior. For orientation, the tubes typically exhibit $U_{0}$ values of about 200 V with gains larger than $5\times 10^{6}$ at 1800 V.
@@
Top: The gain data was derived from the trace charge analysis and the curves are fits for each PMT. The red dotted and blue dash-dotted curves indicate the expected gain scaling for divider A and divider B, respectively, whose data was taken from the 9821B datasheet [9821B_datasheet]. Bottom: The TTS decreases approximately linearly with voltage and lines are linear fits to each tube.
@@ 5.3 Timing
The TTS decreases with voltage across the covered range for all tested tubes. Figure 8 (bottom) shows the individual FWHM of the TTS together with linear fits illustrating this trend. The measured values cluster around the $\sim 2.2$ ns quoted in the 9821B datasheet [9821B_datasheet], but with clear device-specific variations. The TTS and its uncertainty arise from the sampling rate, the constant-fraction timing algorithm, the KDE bandwidth and interpolation procedure for determining the FWHM.
As detailed in [PhD_MRStock2025], introducing an aperture in front of the photocathode reduces TTS by suppressing the spread of photoelectron trajectories. For detailed studies of the prepulse/late pulse components, it is recommended to record datasets containing several millions of SPE events.
@@ 5.4 Temperature and cabling
Temperature scans from $50^{\circ}$C to $20^{\circ}$C were performed using long signal cables for several 9821(Q)B tubes and the R9980. Although the minimum recommended operating temperature for these models is $30^{\circ}$[9821B_datasheet, R9980_datasheet], all tubes remained operational down to $50^{\circ}$C.
The gain increases systematically toward lower temperature by a few tenths of a percent per degree Celsius. This behavior is also tube-dependent and can be explained by changes of the surface properties of the dynodes, affecting the secondary emission factor and thus leading to higher gains [LeoBook]. Additionally, decreased resistance of the PMT divider circuit and signal cables results in higher charge output [QualiPMTs_XENON_2017]. Figure 9 shows the R9980 gain derived from pulse integrals as a function of temperature and voltage. Within uncertainties no robust temperature dependence is observed for the TTS, as illustrated in figure 10. All measured TTS values are compatible with the 0.2 ns specification for the R9980 [R9980_datasheet]. The P/V ratio shows no clear temperature trend. A systematic dependence on cable length is observed. Shorter cables yield slightly higher gain and a smaller apparent TTS due to reduced attenuation and dispersion. This effect exceeds the thermal influence and must be accounted for when comparing different setups.

### v2

@@ 4.3 Transit time spectrum
A KDE with appropriate bandwidth is used to smooth the distribution and to locate minima, which serve to separate pulse populations. Following [25, 10], three populations are identified: prepulses (PPs), main pulses, and late pulses (LPs). Their fractions to the complete data set are counted in the KDE-defined intervals, with an optional dark rate correction estimated from the constant pre-PP region if present. Local full width at half maxima (FWHMs) and full width at quarter maxima (FWQMs) of the pulse populations are taken from the KDE by linear interpolation, and uncertainties are propagated. The three populations are described in the following:
@@
    Prepulses occur $\Delta t_{\text{pre-main}}\simeq 12.6$ ns before the main peak and contribute $0.02\,\%$ of events. The $\text{FWHM}_{\text{PP}}=(1.9\pm 0.1)$ ns is comparable to the main component’s $\text{FWHM}_{\text{TTS}}=(2.0\pm 0.1)$ ns. PPs arise when photons pass the photocathode and strike the first dynode directly [4], yielding smaller charge because the first amplification stage is bypassed [10]. Following [25], adapting a photocathode-to-first-dynode distance of $\sim 2.5$ cm (time-of-flight $t_{\text{ToF}}\simeq 0.1$ ns) and the measured pre-main delay $\Delta t_{\text{pre-main}}$, the PE drift time for the 9821B is $t_{\text{PE\,drift}}=\Delta t_{\text{pre-main}}+t_{\text{ToF}}\simeq 12.7$ ns.
@@
    Main pulses are dominant ($98.58\,\%$) and occur from 8.5-8.5 ns to $11.9$ ns with $\text{FWHM}_{\text{TTS}}=(2.0\pm 0.1)$ ns, which is close to the value in the 9821B datasheet [8]. The TTS reflects variations in PE paths/velocities and secondary-electron production, generally increasing with PMT size. Within the main component we distinguish (i) early pulses from scattering that bypasses an amplification stage (slightly lower charge), (ii) regular pulses that mainly define the TTS, and (iii) delayed pulses from weaker acceleration near photocathode edges [10].
@@
    Late pulses comprise $1.40\,\%$ of events, start at $11.9$ ns and peak at $27.7$ ns. They originate from (in)elastic backscattering of photoelectrons at the first dynode. The longest delays are $\approx 2\,t_{\text{PE\,drift}}\simeq 25.4$ ns, consistent with but slightly below the observed LP peak-to-main pulse peak separation due to trajectory spread [25, 10, 17]. Elastic backscatter tends to cluster near the LP peak with charge similar to normal events, while inelastic cases arrive earlier with reduced charge [4]. A secondary bump broadens the LP component, giving $\text{FWHM}_{\text{LP}}=(5.4\pm 0.1)$ ns and $\text{FWQM}_{\text{LP}}=(14.5\pm 0.1)$ ns.
Figure 7 presents charge-time correlations using the trace integral (top) and pulse integral (bottom) methods. The trace integral method illustrates that prepulses are absent within the SPE interval. Due to the systematically smaller charge of PPs, integrating the full waveform, i.e., signal plus surrounding noise, masks their presence.
@@ 5 Results and discussion
This section summarizes the performance of eight 9821B and two 9821QB PMTs from ET Enterprises as well as one Hamamatsu R9980 PMT assembly, characterized at SPE intensity over a range of supply voltages. The “Q” designation denotes a version of the PMT equipped with a quartz glass window rather than a borosilicate glass window [9]. Although the R9980 PMT requires a negative bias, all voltages are quoted by magnitude. All investigated 9821B devices are equipped with B19A basic hardpin bases, MS75/90 mu-metal shields and E638PFP-01 passive encapsulated voltage dividers. Unless stated otherwise, measurements refer to room temperature and short signal cables in the light‑tight enclosure. Additional details are available in [22].
@@ 5.1 Gain
Figure 8 (top) shows the gain-voltage curves obtained from the trace charge analysis. For each device, an exponential model, $G(U)=G_{0}\exp{\left(U/U_{0}\right)}$, where $G_{0}$ and $U_{0}$ are two fit parameters, describes the data well. Best‑fit curves follow the general scaling suggested by the manufacturer’s divider A benchmark. Dividers A and B are alternative voltage divider configurations provided by the manufacturer for the same PMT type. Divider B employs a modified resistor distribution that supports higher pulsed anode currents and improved linearity compared to divider A [8]. The parameters $G_{0}$ and $U_{0}$ vary tube‑to‑tube, emphasizing device‑specific behavior. For orientation, the tubes typically exhibit $U_{0}$ values of about 200 V with gains larger than $5\times 10^{6}$ at 1800 V.
@@
Top: The gain data was derived from the trace charge analysis and the curves are fits for each PMT. The red dotted and blue dash-dotted curves indicate the expected gain scaling for divider A and divider B, respectively, whose data was taken from the 9821B datasheet [8]. Bottom: The TTS decreases approximately linearly with voltage and lines are linear fits to each tube.
@@ 5.3 Timing
The TTS decreases with voltage across the covered range for all tested tubes. Figure 8 (bottom) shows the individual FWHM of the TTS together with linear fits illustrating this trend. The measured values cluster around the $\sim 2.2$ ns quoted in the 9821B datasheet [8], but with clear device-specific variations. The TTS and its uncertainty arise from the sampling rate, the constant-fraction timing algorithm, the KDE bandwidth and interpolation procedure for determining the FWHM.
As detailed in [22], introducing an aperture in front of the photocathode reduces TTS by suppressing the spread of photoelectron trajectories. For detailed studies of the prepulse/late pulse components, it is recommended to record datasets containing several millions of SPE events.
@@ 5.4 Temperature and cabling
Temperature scans from $50^{\circ}$C to $20^{\circ}$C were performed using long signal cables for several 9821(Q)B tubes and the R9980. Although the minimum recommended operating temperature for these models is $30^{\circ}$[8, 16], all tubes remained operational down to $50^{\circ}$C.
The gain increases systematically toward lower temperature by a few tenths of a percent per degree Celsius. This behavior is also tube-dependent and can be explained by changes of the surface properties of the dynodes, affecting the secondary emission factor and thus leading to higher gains [31]. Additionally, decreased resistance of the PMT divider circuit and signal cables results in higher charge output [26]. Figure 9 shows the R9980 gain derived from pulse integrals as a function of temperature and voltage. Within uncertainties no robust temperature dependence is observed for the TTS, as illustrated in figure 10. All measured TTS values are compatible with the 0.2 ns specification for the R9980 [16]. The P/V ratio shows no clear temperature trend. A systematic dependence on cable length is observed. Shorter cables yield slightly higher gain and a smaller apparent TTS due to reduced attenuation and dispersion. This effect exceeds the thermal influence and must be accounted for when comparing different setups.